# Advances in Protein Representation Learning: Methods, Applications, and Future Directions

### v1

@@ Abstract
Protein representation learning (PRL) has emerged as a transformative approach to advancing our understanding of proteins and addressing critical challenges in molecular biology, biotechnology, and medicine. This review provides a comprehensive overview of PRL methodologies, categorized into five major approaches: feature-based, sequence-based, structure-based, multimodal, and complex-based. By analyzing the strengths and limitations of each approach, we offer insights to guide the selection of appropriate PRL strategies based on specific biological and computational requirements, ensuring their effective application across diverse research domains.
@@
Beyond methodological advancements, we highlight key databases that underpin PRL research and discuss its diverse applications, including protein property prediction, structure modeling, protein design, and drug discovery. Despite the substantial progress in PRL, several open challenges remain, presenting opportunities for future research. Key directions for advancement include extending PRL methodologies to genomic representation learning, enhancing model scalability and generalization, and improving interpretability to foster broader accessibility and adoption. Addressing these challenges will be crucial in refining PRL methodologies and expanding their impact across biological and biomedical research.

### v2

@@ 10. Conclusion
Protein representation learning (PRL) has emerged as a transformative approach to advance our understanding of proteins and to address critical challenges in molecular biology, biotechnology, and medicine. This review provides a comprehensive overview of PRL methodologies, categorized into five main approaches: feature-based, sequence-based, structure-based, multimodal, and complex-based. By analyzing the strengths and limitations of each approach, we offer insights to guide the selection of appropriate PRL strategies based on specific biological and computational requirements, ensuring their effective application across diverse research domains.
@@
Beyond methodological advancements, we highlight key databases that underpin PRL research and discuss its diverse applications, including protein property prediction, structure modeling, protein design, and drug discovery. Despite substantial progress in PRL, several open challenges remain, which present opportunities for future research. Key directions for advancement include extending PRL methodologies to genomic representation learning, enhancing model scalability and generalization, and improving interpretability to foster broader accessibility and adoption. Addressing these challenges will be crucial to refining PRL methodologies and expanding their impact on biological and biomedical research.